OpenStudy (anonymous):

The power, p, produced by an electric circuit varies directly as the square of the voltage, V, and inversely as the resistance, R. Write an equation that matches that description.

OpenStudy (anonymous):

p=(v^2)/r

OpenStudy (anonymous):

hmmm the only answer in the multiple choices that is similar is: p=kR/v^2

OpenStudy (anonymous):

The question is worded strangly, it could mean that p^2=v, in which case p=k(squrt(v))/r. I did forget the constant k in my initial responce, but I am certain, by the definition of inverse relations that p=x/r, where x is a function of v.

OpenStudy (anonymous):

So.... is that wrong? there is: A. p= (k)/(v^2r) B. p= (kr)/(v^2) C. p= (kv^2)/(r) D. p= kv^2r

OpenStudy (anonymous):

C. it varies directly with the square of voltage, hence the V^2, and inversely with resistance hence the 1/R. the k is just an arbitrary constant that you have to tag along as a scalar hence, p = (kV^2)/R

OpenStudy (anonymous):

thank you!!!